OpenStudy (anonymous):

An airplane travels 900 miles from Houston to Miami in 6 hours against the wind. On its return trip, with the wind, it only takes 5 hours. Find the rate of the airplane with no wind. Find the rate of the wind.

OpenStudy (anonymous):

Let the ground speed be x And the wind speed be y AGAINST WIND Distance = 900 miles. Distance = Rate * time 900= (x-y)*6 WITH WIND Distance = 900 miles. Distance = Rate * time 900= (x+y)*5 6x-6y=900 x-y= 150 5x+5y=900 x+y=180 x-y= 150 x+y=180 2x=330 x= 165 mph = ground speed ( With no wind ) y= 15 The ground speed of the plane is 165 mph and wind speed = 15 mph

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem :D

OpenStudy (anonymous):

How Did You get x-y= 150 ?

OpenStudy (anonymous):

that's just how you find speed without wind resistance. 900 = (x-y) *6 900/6 = x-y 150 = x-y

OpenStudy (anonymous):

How Did You Get 2x=330 ?

OpenStudy (anonymous):

Never mind Thanks

OpenStudy (anonymous):

so what was the answer?

OpenStudy (anonymous):

x=165 mph y=15 Wind Speed

OpenStudy (anonymous):

I'm sorry I didn't answer I'm doing my homework

OpenStudy (anonymous):

Its Okay I Believe I Know How To Do This Now :D

OpenStudy (anonymous):

Glad I helped :D